# Beyond Statistical Estimation: Differentially Private Individual Computation in the Shuffle Model

### v1

@@ III-A Privacy Definitions
The data processing inequality is considered a key feature of distance measures used for evaluating privacy. It asserts that the privacy guarantee cannot be weakened by further analysis of a private mechanism’s output.
@@ III-B The Classical Shuffle Model
III-C Security Primitives
We employ several well-established security primitives for secure point-to-point communication, including key aggreement, symmetry encryption and digital signature.
A Diffie-Hellman key agreement protocol with security parameter $\lambda$ comprises two primary steps:
  • $sk_{i},pk_{i}\leftarrow\textsf{DH-Gen}(\lambda)$: a party $i$ generates a public-private key pair $(sk_{i},pk_{i})$, where $pk_{i}$ is the public key and $sk_{i}$ is the secret/private key;
  • $K_{ij}\leftarrow\textsf{DH-Compute}(sk_{i},pk_{j})$: the two parties involved in the key exchange use their private keys $sk_{i}$ and $sk_{j}$ to compute a shared key $K_{ij}$ from the public keys $pk_{j}$ and $pk_{i}$ of each party (i.e., $K_{ij}=K_{ji}$.
The resulting shared secret key, $K_{ij}$, can be used for subsequent data symmetry encryption/decryption. Specifically, a symmetry encryption scheme with secret key $sk$ consists of two steps:
  • $c\leftarrow\textsf{Enc}(sk,x)$: generates a ciphertext $c$ by encrypting data $x$ with the encryption key $sk$;
  • $x\leftarrow\textsf{Dec}(sk,c)$: decrypts ciphertext $c$ with the key $sk$, outputting $x$.
A digital signature scheme permits the receipt to authorize ciphertexts and to ensure message integrity, through the following steps:
  • $t\leftarrow\textsf{Sign}(sk,c)$: returns a tag $t$ as the output of signing data $c$ with the private key $sk$;
  • $accept/reject\leftarrow\textsf{Verify}(pub,t,c)$: outputs $accept$ when $t=\textsf{Sign}(sk,c)$, and outputs $reject$ with a probability of $1-negl(k)$ when $t\neq\textsf{Sign}(sk,c)$.
@@ IV Problem Settings
We aim to extend the capability of the shuffle model to support private individual computation. This would have multiple benefits: (1) we can get a formal privacy guarantee (in the DP sense) on a much wider range of data-driven computational tasks; (2) because the computation is done on already sanitized data, we can avoid heavy cryptographic protocols and scale to large data and/or user size; (3) the privacy amplification effect will allow much better utility, compared to sanitizing and computing directly using LDP. In this section, we will set the stage by first presenting a few motivating applications, and then formally defining private individual computation as an ideal functionality.
Figure 1: An illustration of conceptual overview of PIC model for spatial crowdsourcing.
@@ IV-A Motivating Applications
We describe three prevalent exemplar computation tasks: spatial crowdsourcing, location-based services, and federated learning with incentives:
Spatial crowdsourcing. A spatial crowdsourcing system normally consists of three roles: users, workers, and the orchestrating server (the service platform), and proceed as following major steps:
@@
    Task submission: each participated user $i\in G_{a}$ submits a task (e.g., taxi calling requests, sensing/photo requests) $x_{i}=(i,l_{i},v_{i})$ containing the location $l_{i}$ and possibly other information $v_{i}$;
@@
    Task performing: users and workers retrieve matching results $M$ and collaboratively complete the task.
Location-based services. In a location-based service, a set of users and a server interact as the following:
@@
Federated learning with incentives. Federated learning involves a set of clients and a server:
@@
    Submitting gradient: Each client $i\in G_{a}$ in each epoch computes an intermediate gradient information $x_{i}\in[-1,1]^{d}$ with its local model and data, then submits to the server;
@@
    Computing incentive: The server computes the averaged gradient $\overline{x}=\frac{1}{n}\sum_{i\in[n]}x_{i}$. To incentive participation, the server may reward clients with monetary tokens (e.g., via cryptocurrency) according to a profit allocation algorithm $V:[-1,1]^{d\times n}\times[-1,1]^{d}\mapsto[0,1]$ (e.g., Shapley value [69]).
@@
    Receiving incentive: Each client retrieves the token and claims its monetary incentive.
A fundamental distinction between the above applications and those currently studied in the shuffle model is that now each participant expects an output that differs individually, instead of a single collectively aggregated output. Informally, in such applications, we would need to safeguard data privacy so that apart from the party who generates the data, no one should be certain about that party’s data (up to the leakage allowed by DP). To be able to amplify privacy, we also need to maintain anonymity so that for any message, an adversary (e.g. the server, an observer, an unmatched user) should only know this message comes from a user belonging to a particular group, but nothing more. Note that following the convention in the shuffle model, the shuffler is trusted for providing anonymity. So the shuffler will know the random permutation used in the shuffling process and will not leak it to other parties, although we prevent the shuffler from observing plaintext messages through encryption.
@@ IV-B The Ideal Functionality
Functionality $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$ Parameters: $m\in\mathbb{N}$; $m$ groups of parties $G_{1},G_{2},\cdots,G_{m}$, where $n_{i}=|G_{i}|$ denotes the number of parties in group $G_{i}$; the data randomization mechanisms $\mathcal{R}_{i}$ for group $G_{i}$; a server ${S}$. Functionality: Upon receiving $n=\sum_{i\in[m]}n_{i}$ inputs $\{x_{i,j}\}_{i\in[m],j\in[|G_{i}|]}$ from all clients, and the description of a function $f$ to be computed over parties’ inputs from the server, do the following: Compute $x^{\prime}_{i,j}\leftarrow R_{i}(x_{i,j})$ for all $i\in[m]$ and $j\in[|G_{i}|]$. Sample $m$ random permutations $\pi_{1},\pi_{2},\cdots,\pi_{m}$, where $\pi_{i}:[n_{i}]\rightarrow[n_{i}]$, perform shuffle over inputs and obtain $L=\left(\{x_{1,\pi_{1}(j)}^{\prime}\}_{j\in[n_{1}]},\cdots,x_{m,\pi_{m}(j)}^{\prime}\}_{j\in[n_{m}]}\right)$, and compute ($\left(\{y_{1,\pi_{1}(j)}\}_{j\in[n_{1}]},\cdots,\{y_{m,\pi_{m}(j)}\}_{j\in[n_{m}]}\right)\leftarrow f(L)$. Send $y_{i,\pi_{i}(j)}$ to party $u_{i,j}$ for all $i\in[m]$ and $j\in n_{i}$. Additionally, send $(L,f(L))$ to $S$.
@@
Essentially, the ideal functionality represents a fully trusted party that interacts with $m$ groups of clients and one server $S$. It sanitizes the input from each client, then shuffles them randomly, and applies a function from the server to compute the output for each client. At the end, each client receives its function output, and the server gets the list of shuffled sanitized data from the clients and a list of the function output. This captures the functional requirement of individual computation in the real world: a computing task is performed by a server using the joint inputs from a set of clients (which are sanitized and shuffled), and each client gets an individualized output. It also captures the privacy requirements: each party should receive precisely the specified output, nothing more.
Remark A careful reader may notice that the ideal function does not capture differential privacy. This is intentional for the sake of security analysis. Later in the security analysis, we will decouple different privacy requirements into two sets of proofs: the first set shows that our concrete protocol executed by real-world parties realizes the ideal functionality, meaning that no additional information about parties’ inputs is revealed, except the output given to each party; then we prove that the output conforms with differential privacy.
@@ V-A The Protocol
We now present a concrete protocol for private individual computation. Because there are already secure protocols for shuffling, to simplify the description, we present the protocol in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model, in which parties can communicate as usual, and in addition have access to an ideal functionality $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ that does the shuffling. The ideal functionality $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ (Figure 3) is parameterized with a list of parties that are corrupted by the adversary and collude with the server. For those parties, the adversary should know the correspondence between their messages before and after shuffling, hence $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ leaks this part of the permutation to the adversary.
@@
The Private Individual Computation protocol is outlined below:
@@
    We denote the $j$-th client in group $G_{i}$ as $u_{i,j}$. Each client generates a key pair $(pk_{i,j},sk_{i,j})\leftarrow\textsf{Gen}(\lambda)$. Each client then randomizes its private data $x_{i,j}$ with mechanism $\mathcal{R}_{i}$ and obtains $x_{i,j}^{\prime}\leftarrow\mathcal{R}_{i}(x_{i,j})$. Then the sanitized input is concatenated with its own public key, and encrypted with the server’s public key $x_{i,j}^{\prime\prime}\leftarrow\textsf{Enc}_{pk_{c}}(pk_{i,j}||x^{\prime}_{i,j})$.
@@
    Each client in group $G_{i}$ invokes $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ with $x_{i,j}^{\prime\prime}$, and $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ outputs the shuffled messages $\{x_{i,\pi_{i}(j)}^{\prime\prime}\}_{j\in[n_{i}]}\leftarrow\mathcal{S}(\{x_{i,j}^{\prime\prime}\}_{j\in[n_{i}]})$ to the server, where $\pi^{-1}_{i}$ is the (secret) random permutation used during $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ for group $G_{i}$.
@@
    It then computes the function $f$ over $L$ to produce output for each anonymous client:
    ($\left(\{y_{1,\pi_{1}(j)}\}_{j\in[n_{1}]},\cdots,\{y_{m,\pi_{m}(j)}\}_{j\in[n_{m}]}\right)\leftarrow f(L)$
@@
    Every client downloads the list, finds in the list the entry with its own public key, and decrypts the payload to get the computation result.
@@ V-B Security Analysis
We first show that our protocol securely realizes the ideal functionality $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$ in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model. This means the protocol leaks no more information than what is allowed by $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ and $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$. More precisely, each client gets their output from $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$, the server gets $(L,f(L))$, and in the case of colluding with some clients in $G_{i}$, the partial permutation $\mathcal{L}(\pi_{i})$. Formally we have the following theorem:
@@
The proof can be found in Appendix -J. The proof is simulation-based. It shows that for any PPT adversary $\mathcal{A}$ in the real world, there exists a PPT simulator $\mathcal{S}$ in the ideal world that can generate a simulated view given the corrupted parties’ inputs and outputs. Security means that the simulated view is indistinguishable from the view of real-world execution.
@@
In particular, when all users in $G_{i^{*}}$ use an identical $\epsilon$-LDP mechanism as the data randomizer $\mathcal{R}_{i^{*}}$, and for $n^{\prime}_{i^{*}}=|G_{i^{*}}-C_{i^{*}}|\geq\tilde{\Omega}(e^{\epsilon})$ we have:
$\small\epsilon_{c}=\log\Bigg{(}1+\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\Big{(}\sqrt{\frac{32(e^{\epsilon}+1)\log 4/\delta}{n^{\prime}_{i^{*}}}}+\frac{4(e^{\epsilon}+1)}{n^{\prime}_{i^{*}}}\Big{)}\Bigg{)}.\normalsize$ (1)
The analysis can be divided into two cases: in the first case, $S\in\mathcal{C}$, i.e. S is corrupted. In this case, since the honest client locally randomizes their input, the input enjoys at least $\epsilon$-DP. Then the shuffling will amplify the privacy guarantee. Since the corrupted server receives the partial permutation as the leakage, the amplification depends on the number of uncorrupted users in the same group ($n^{\prime}_{i^{*}}$) as the honest user. The amplified $\epsilon_{c}$ can then be derived following [31, 32]. In the second case where $S\not\in\mathcal{C}$, the knowledge of the adversary is $y_{i,j}$ for each corrupted user $u_{i,j}\in\mathcal{C}$. The tricky part is that how much information $y_{i,j}$ leaks depends on the function $f$ being evaluated by the server. Hence we consider the worst case where $f$ output $y_{i,j}=L$. Continuing along the same line of thought, we conclude that the level of differential privacy assurance is no less than in the first case. The full proof can be found in Appendix -A.

### v2

@@ III-A Privacy Definitions
The data processing inequality is considered a key feature of distance measures (e.g., Hockey-stick divergence) used for evaluating privacy. It asserts that the privacy guarantee cannot be weakened by further analysis of a mechanism’s output.
@@ III-B The Classical Shuffle Model
III-C Public Key Encryption
[…]
  • Gen: takes as input the security parameter $\lambda$ and outputs a pair of keys $(pk,sk)$, where $pk$ denotes the public key and $sk$ the private key.
  • Enc: takes as input a public key $pk$ and a message $m$ from the plaintext space. It outputs a ciphertext $c\rightarrow\textsf{Enc}_{pk}(m)$.
  • Dec: takes as input a private key $sk$ and a ciphertext $c$, and outputs a message $m$ or a special symbol $\bot$ denoting failure.
It is required that for every $(pk,sk)$ and plaintext message $m$, it holds that $\textsf{Dec}_{sk}(\textsf{Enc}_{pk}(m))=m$. The IND-CPA security ensures the scheme leaks no useful information under a chosen plaintext attack.
@@ IV Problem Settings
We now introduce Private Individual Computation, a new paradigm for privacy-preserving computation that offers several benefits: (1) It provides a formal privacy guarantee (in the differential privacy sense) for a wide range of data-driven computational tasks. (2) Since the computation is performed on user-sanitized data, it avoids the need for heavy cryptographic protocols, allowing for scalability to handle large datasets and user bases. (3) The privacy amplification effect results in significantly better utility compared to direct sanitization and computation using Local Differential Privacy (LDP). In this section, we will first present a few motivating applications and then formally define Private Individual Computation as an ideal functionality.
Figure 1: An illustration of taxi-hailing in the PIC model. Besides (sanitized) location information, each user also encapsulates a one-time random public key into the message to the shuffler.
@@ IV-A Motivating Applications
We describe three prevalent exemplar computation tasks: spatial crowdsourcing, location-based social systems, and federated learning with incentives:
Spatial crowdsourcing. A spatial crowdsourcing system typically consists of three roles: users, workers, and the orchestrating server (the service platform). It proceeds through the following major steps:
@@
    Task submission: each user $i\in G_{a}$ submits a task (e.g., taxi calling requests, sensing/photo requests) $x_{i}=(i,l_{i},v_{i})$ containing the location $l_{i}$ and possibly other information $v_{i}$;
@@
    Task performing: users and workers retrieve matching results and collaboratively complete the task.
One example of taxi-haling is illustrated in Figure 1.
Location-based social systems. In a location-based social system, a set of users and a server interact as the following:
@@
Federated learning with incentives. Federated learning involves a set of users and a server:
@@
    Submitting gradient: Each user $i\in G_{a}$ in each epoch computes an intermediate gradient information $x_{i}\in[-1,1]^{d}$ with its local model and data, then submits to the server;
@@
    Computing incentive: The server computes the average gradient $\overline{x}=\frac{1}{n}\sum_{i\in[n]}x_{i}$. To incentive participation, the server may reward users with monetary tokens (e.g., via cryptocurrency) according to a profit allocation algorithm $V:[-1,1]^{d\times n}\times[-1,1]^{d}\mapsto[0,1]$ (e.g., Shapley value [71]).
@@
    Receiving incentive: Each user retrieves the token and claims its monetary incentive.
A fundamental distinction between the above applications and those currently studied in the shuffle model is that each participant now expects an output that differs individually, rather than a single collectively aggregated output. Another distinction is that a participant might need to securely communicate with other participants after receiving the individualized computation results, such as the matched user and worker in spatial crowdsourcing will communicate with each other to accomplish the task, and the matched users in location-based social systems would like to securely contact each other afterward. Informally, in such applications, it is necessary to safeguard data privacy so that, aside from the party who generates the data, no one can be certain about that party’s data (up to the leakage allowed by differential privacy). To amplify privacy, we also need to maintain anonymity so that for any given message, an adversary (such as the server, an observer, or an unmatched user) should only know that this message comes from a user within a particular group, but nothing more. Following the convention in the shuffle model, the shuffler is trusted to provide anonymity. The shuffler knows the random permutation used in the shuffling process and will not leak it to other parties, although the shuffler is prevented from observing plaintext messages through encryption.
@@ IV-B The Ideal Functionality
Functionality $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$ Parameters: $m\in\mathbb{N}$; $m$ groups of parties $G_{1},G_{2},\cdots,G_{m}$, where $n_{i}=|G_{i}|$ denotes the number of parties in group $G_{i}$; the data randomization mechanisms $\mathcal{R}_{i}$ for group $G_{i}$; a server ${S}$. Functionality: Upon receiving $n=\sum_{i\in[m]}n_{i}$ inputs $\{x_{i,j}\}_{i\in[m],j\in[|G_{i}|]}$ from all users, and the description of a function $f$ to be computed over parties’ inputs from the server, do the following: Compute $x^{\prime}_{i,j}\leftarrow R_{i}(x_{i,j})$ for all $i\in[m]$ and $j\in[|G_{i}|]$. Sample $m$ random permutations $\pi_{1},\pi_{2},\cdots,\pi_{m}$, where $\pi_{i}:[n_{i}]\rightarrow[n_{i}]$, perform shuffle over inputs and obtain $L=\left(\{x_{1,\pi_{1}(j)}^{\prime}\}_{j\in[n_{1}]},\cdots,\{x_{m,\pi_{m}(j)}^{\prime}\}_{j\in[n_{m}]}\right)$, and compute ($\left(\{y_{1,\pi_{1}(j)}\}_{j\in[n_{1}]},\cdots,\{y_{m,\pi_{m}(j)}\}_{j\in[n_{m}]}\right)\leftarrow f(L)$. Send $y_{i,\pi_{i}(j)}$ to party $u_{i,j}$ for all $i\in[m]$ and $j\in n_{i}$. Additionally, send $(L,f(L))$ to $S$.
@@
Essentially, the ideal functionality represents a fully trusted party that interacts with $m$ groups of users and one server $S$. It sanitizes the input from each user, shuffles the inputs randomly, and then applies a function to compute the output for each user. In the end, each user receives their individual function output, while the server receives a list of the shuffled sanitized user data and a list of the function outputs. It captures the functional requirements of individual computation in the real world: a server performs a computing task using the joint inputs from a set of users (which are sanitized and shuffled), and each user receives an individualized output. It also captures the security requirements: each party receives precisely the specified output, and nothing more.
Remark 1 A careful reader may notice that the ideal functionality does not explicitly capture differential privacy. This omission is intentional for the sake of security analysis. In the security analysis, we decouple the privacy requirements into two sets of proofs: the first set demonstrates that our concrete protocol, when executed by real-world parties, realizes the ideal functionality. This means no additional information about parties’ inputs is revealed, except for the output given to each party. The second set of proofs establishes that the output given to each party conforms with differential privacy.
@@ V-A The Protocol
We now present a concrete protocol for PIC. Because there are already secure protocols for shuffling, to simplify the description, we present the protocol in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model, in which parties can communicate as usual, and in addition have access to an ideal functionality $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ that does the shuffling. The ideal functionality $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ (Figure 3) is parameterized with a list of parties that are corrupted by the adversary and collude with the server. For those parties, the adversary should know the correspondence between their messages before and after shuffling, hence $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ leaks this part of the permutation to the adversary.
@@
The PIC protocol is outlined below:
@@
    We denote the $j$-th user in group $G_{i}$ as $u_{i,j}$. Each user generates a key pair $(pk_{i,j},sk_{i,j})\leftarrow\textsf{Gen}(\lambda)$. Each user then randomizes their private data $x_{i,j}$ with mechanism $\mathcal{R}_{i}$ and obtains $x_{i,j}^{\prime}\leftarrow\mathcal{R}_{i}(x_{i,j})$. Then the sanitized input is concatenated with their own public key, and encrypted with the server’s public key $x_{i,j}^{\prime\prime}\leftarrow\textsf{Enc}_{pk_{c}}(pk_{i,j}||x^{\prime}_{i,j})$.
@@
    Each user in group $G_{i}$ invokes $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ with $x_{i,j}^{\prime\prime}$, and $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ outputs the shuffled messages $\{x_{i,\pi_{i}(j)}^{\prime\prime}\}_{j\in[n_{i}]}\leftarrow\mathcal{S}(\{x_{i,j}^{\prime\prime}\}_{j\in[n_{i}]})$ to the server, where $\pi^{-1}_{i}$ is the (secret) random permutation used during $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ for group $G_{i}$.
@@
    It then computes the function $f$ over $L$ to produce output for each anonymous user:
    ($\left(\{y_{1,\pi_{1}(j)}\}_{j\in[n_{1}]},\cdots,\{y_{m,\pi_{m}(j)}\}_{j\in[n_{m}]}\right)\leftarrow f(L).$
@@
    Every user downloads the list, finds in the list the entry with their own public key, and decrypts the payload to get the computation result.
Remark 2 In the final step, we adopt the simplest strategy for users to retrieve their results without the server knowing which entry belongs to whom. If bandwidth is a concern, it can be replaced by a more sophisticated (and computationally more expensive) Private Information Retrieval protocol (e.g. [46]).
Remark 3 To eventually accomplish the spatial crowdsourcing (e.g., taxi-hailing services) or location-based social system tasks, the computation result $y_{i,\pi_{i}(j)}$ will contain the matched users of the user $\pi_{i}(j)\in G_{i}$. That is, $y_{i,\pi_{i}(j)}$ will encapsulate a list of public keys and noisy location information about the matched users of the user $\pi_{i}(j)\in G_{i}$. After receiving the individualized computation results in the final step, each user $\pi_{i}(j)$ can then securely contact the matched users using the public keys in $y_{i,\pi_{i}(j)}$ (possibly with the help of a public communication channel, such as a public bulletin).
@@ V-B Security Analysis
We first show that our protocol securely realizes the ideal functionality $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$ in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model. This means the protocol leaks no more information than what is allowed by $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ and $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$. More precisely, each user gets their output from $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$, the server gets $(L,f(L))$, and in the case of colluding with some users in $G_{i}$, the partial permutation $\mathcal{L}(\pi_{i})$. Formally we have the following theorem:
@@
The proof can be found in Appendix -A. The proof is simulation-based. It shows that for any PPT adversary $\mathcal{A}$ in the real world, there exists a PPT simulator $\mathcal{S}$ in the ideal world that can generate a simulated view given the corrupted parties’ inputs and outputs. Security means that the simulated view is indistinguishable from the view of real-world execution.
@@
In particular, when all users in $G_{i^{*}}$ use an identical $\epsilon$-LDP mechanism as the data randomizer $\mathcal{R}_{i^{*}}$, and for $n^{\prime}_{i^{*}}=|G_{i^{*}}-C_{i^{*}}|\geq 8(e^{\epsilon}+1)\log(2/\delta)$ we have:
$\epsilon_{c}=\log\Bigg{(}1+\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\Big{(}\sqrt{\frac{32(e^{\epsilon}+1)\log 4/\delta}{n^{\prime}_{i^{*}}}}+\frac{4(e^{\epsilon}+1)}{n^{\prime}_{i^{*}}}\Big{)}\Bigg{)}.$ (1)
The analysis can be divided into two cases: in the first case, $S\in\mathcal{C}$, i.e. S is corrupted. In this case, since the honest user locally randomizes their input, the input enjoys at least $\epsilon$-DP. Then the shuffling will amplify the privacy guarantee. Since the corrupted server receives the partial permutation as the leakage, the amplification depends on the number of uncorrupted users in the same group ($n^{\prime}_{i^{*}}$) as the honest user. The amplified $\epsilon_{c}$ can then be derived following [34, 35]. In the second case where $S\not\in\mathcal{C}$, the knowledge of the adversary is $y_{i,j}$ for each corrupted user $u_{i,j}\in\mathcal{C}$. The tricky part is that how much information $y_{i,j}$ leaks depends on the function $f$ being evaluated by the server. Hence we consider the worst case where $f$ output $y_{i,j}=L$. Continuing along the same line of thought, we conclude that the level of differential privacy assurance is no less than in the first case. The full proof can be found in Appendix -B.
Remark 4 After decryption, the server obtains $pk_{i,j}||x^{\prime}_{i,j}$, where $pk_{i,j}$ is a public key not sanitized by the local randomizer. Despite the presence of the public key, $pk_{i,j}||x^{\prime}_{i,j}$ and $x^{\prime}_{i,j}$ are equivalent in terms of privacy amplification. This is because (1) the public key is random and generated independently of $x_{i,j}$, so prepending it to $x^{\prime}_{i,j}$ does not affect the local DP guarantee—it is the same as $x^{\prime}_{i,j}$ itself; and (2) all public keys follow an identical probability distribution across all users in the group $G_{i}$, ensuring that the privacy amplification effect via shuffling is not degraded.